You win the lottery. You can either have $850,000 right now, or $1,000,000 paid in annual installments of $200,000 starting next year. If the interest rate is 6%(0.06), which will you choose? Why? What if the interest rate is 4%(0.04)? Explain.
Answer #1

When I = 6%

Present value of installment payment = 200000 * (P/A, 6%,5)

= 200000 * 4.212364

= 842472.76

Present value of second option = 850000

As the present value of getting payment now is more, it should be selected

When I = 4%

Present value of installment payment = 200000 * (P/A, 4%,5)

= 200000 * 4.451822

= 890364.47

Present value of second option = 850000

As the present value of getting payment in installments is more, it should be selected